INTENSITIES, COLLISION-BROADENED HALF-WIDTHS, AND COLLISION-INDUCED LINE SHIFTS IN THE FUNDAMENTAL AND OVERTONE BANDS OF ^{12}C^{16}O RELEVANT TO THE TERRESTRIAL AND JOVIAN ATMOSPHERES

\underline{KEEYOON SUNG} and PRASAD VARANASI, Institute for Terrestrial and Planetary Atmospheres, State University of New York at Stony Brook, Stony Brook, New York 11794-5000.

Absolute intensities, self-broadened, air-broadened and H_{2}-broadened half-widths, and self-induced, air-induced and H_{2}- induced line shifts were measured in the second overtone (3-0) band of ^{12}C^{16}O at 296 K using a Fourier-transform spectrometer and a multi-spectral-line-fitting technique. The absolute intensity of the band was determined to be (4.62 ± 0.06) × 10^{−22} cm · molecule^{−1}. The square of the transition moment, |μ_{0}|^{2}, of this band was deduced to be (1.712 ± 0.004) × 10^{−7}Debye^{2}. The two Herman-Wallis factors are estimated to be C = (1.185 ± 0.009) × 10^{−2} and D = (1.014 ± 0.069) × 10^{−4}. H_{2}-broadened half-widths and H_{2}-induced line shifts in the (1-0) and (2-0) bands of ^{12}C^{16}O were measured at several temperatures between 83 and 302 K relevant to Jovian planets. CO_{2}-broadened half-widths and CO_{2}-induced line shifts in the fundamental band of ^{12}C^{16}O were also measured at 202, 245 and 300 K. Self- and air-broadened line widths decrease with m, where m is −J for P-branch and J+1 for Rbranch. However, H_{2}-broadened line widths, whose values of the lines in the fundamental band were found to be in excellent agreement with those reported by Varanasi et al (1987), exhibit no significant dependence on m for all the bands measured. While the line widths in the CO (3-0) band measured at the room temperature were nearly the same as those in the 1-0 and 2-0 bands in the self- and air-broadening studied, the collision-induced line shifts were found to be much larger in the 3-0 band. This was also true of H_{2}-broadened spectra. Each mean value of n and ξ, the exponents describing temperature dependence, are determined to be 0.65 for H_{2}-broadened line widths and 1.00 for H_{2}-induced-induced line shifts both in the fundamental and the first overtone band.